### v2

@@ Abstract
Biomarkers play a central role in medicine’s gradual progress towards proactive, personalized precision diagnostics and interventions. However, finding biomarkers that provide very early indicators of a change in health status, particularly for multi-factorial diseases, has been challenging. Discovery of such biomarkers stands to benefit significantly from advanced information processing and means to detect complex correlations, which quantum computing offers. In this perspective paper, quantum algorithms, particularly in machine learning, are mapped to key applications in biomarker discovery. The opportunities and challenges associated with the algorithms and applications are discussed. The analysis is structured according to different data types — multi-dimensional, time series, and erroneous data — and covers key data modalities in healthcare — electronic health records (EHRs), omics, and medical images. An outlook is provided concerning open research challenges.
@@ 1 Introduction
A biomarker is a defined characteristic that is measured as an indicator of normal or pathological biological processes, or of response to an exposure or intervention [1]. The study of biomarkers arguably dates back to at least the beginning of the last century with Karl Landsteiner’s discovery of the ABO blood types in 1901 (recognized with the 1930 Nobel Prize in Physiology or Medicine). Biomarkers consist of both biological moleculesand other medical signs that indicate diseases, in contrast to symptoms as well as more general medical endpoints. Biomarkers can be based on histological, molecular, physiological, and radiographic characteristics, and a biomarker is distinct from an endpoint or a clinical outcome. They may also take the form of aggregate high-dimensional features.
@@
For example, clinical trials in neurology and psychiatry are time-intensive, expensive, and subjective, and have lower success rates than other indications [3]. Within neuro-psychiatry, pain represents a large unmet need with respect to the lack of reliable biomarkers. Hence, quantitative and physiological biomarkers are urgently needed for a comprehensive and reliable diagnosis and assessment of pain in healthcare and clinical trials. This is illustrative of the impact better biomarkers can have across medicine.
Quantum computing, on the other hand, is a new form of information processing, including novel hardware and software, that leverages quantum mechanical effects such as quantum entanglement, interference, and superposition. It is the only known form of computing that can provide, at least for certain problems, exponential speedups compared with classical approaches [4]. Hundreds of use cases are meanwhile being explored across industry and academia. Quantum computing is not just a faster way of addressing problems; it represents an entirely different way to find solutions. The benefits also go beyond speed; depending on the application and the use case, the main benefit could be accuracy [5], energy efficiency [6], or the ability to handle difficult datasets (for instance, high-dimensional or highly noisy ones) [7, 8].
The focus of this review is on (gate-based) quantum computing and its potential for computational problems in biomarker discovery. Nevertheless, it is important to note that there are further quantum technologies with medical applications and biomarker relevance, in particular quantum annealingand quantum sensing. In the rest of the paper, quantum computing and relevant quantum algorithms are first summarized in Section 2 before the application of quantum algorithms to biomarker discovery problems is described in Section 3. Open research challenges are discussed in Section 4 and conclusions are drawn in Section 5.
@@ 2 Quantum computing
In quantum computing, algorithms execute operations in Hilbert space. This is a high-dimensional vector space that represents the state of the qubits and is naturally suited to computing outcome probabilities (via an inner product). If the input of a quantum algorithm is classical data that is not represented by quantum states, data embedding is required in order to perform operations on a quantum computer. While there are various ways to embed classical data into quantum states, the number of available qubits restricts the feature space size of the data we can tackle with near-term devices. For instance, angle encoding [9] or instantaneous quantum polynomial encoding [10] require a linear scaling of the number of qubits with respect to the number of features. Alternatively, amplitude encoding provides a logarithmic scaling in the number of qubits [9]. However, without ancillary qubits, amplitude encoding results in deep circuits [11]. In practice, when the feature space exceeds the capacity of available quantum hardware, classical dimensionality reduction algorithms are applied as a pre-processing step in quantum pipelines. Still, with ongoing advancements in quantum computing, biomarker discovery and other quantum applications may eventually be performed without the need for significant (classical) pre-processing.
Quantum circuits are fundamental to the architecture of quantum computation, providing the framework through which quantum algorithms are designed and executed. At a high level, a quantum circuit is a sequence of quantum gates applied to qubits. Each gate performs a specific operation on one or more qubits. Fault-tolerant quantum algorithms typically require running deep quantum circuits on tens of thousands to millions of qubits, making them unsuitable for near-term devices. On the other hand, variational quantum algorithms (VQAs)utilize shallower circuits and can produce meaningful results. This characteristic has made VQAs highly popular for near-term quantum computing applications. VQAs are hybrid quantum-classical algorithms that involve an iterative loop between quantum and classical hardware. First, quantum circuits that have parametrized gates, called the ansätze or parametrized quantum circuits (PQCs), are employed to obtain samples from a quantum computer. Then, a classical computer optimizes the parameters of the quantum circuit with respect to the loss function of the problem. Most quantum machine learning models belong in the class of VQAs and can be evaluated on quantum computers. Alternatively, quantum computers could be leveraged as sub-routines within classical algorithms. One example is the thermal state preparation during Boltzmann machine training. However, the performance of VQAs is highly sensitive to hardware noise, necessitating the use of error mitigation techniques to partially counteract noise-induced errors [12].
@@
Quantum computing is naturally suited for quantum data (which describes quantum states) [13]. However, it also holds great promise in solving problems using classical data too. In particular, there has been much progress in the last few years in understanding how quantum computers can solve computational problems that can be formulated as cryptography, machine learning (ML), optimization, or simulation (of nature) problems. In this paper, we focus on analyzing classical data with the help of quantum algorithms, including multi-dimensional, time series, and erroneous data, which is discussed in detail in Section 3.
@@ 2.1 Quantum machine learning
The majority of QNNs are based on PQCs. The parameters of PQCs are optimized following the VQA paradigm. A typical QNN begins with data embedding, followed by a series of gates. One of the most general QNN models for classification is the variational quantum classifier (VQC), which is essentially a PQC. Another example of a VQC is the circuit-centric quantum classifier, one of the earliest proposed generic QNNs. The data re-uploading classifier is a notable QNN architecture that was shown to have universal approximation properties [15]. The data re-uploading classifier features a processing layer composed of data encoding and unitary operations with trainable parameters. This layer is repeated a number of times, each layer having independent parameters.
@@
Dimensionality reduction Covariates for biomarker trait regression models QPCA, QLDA, QSFA, QIsomap
Classification Early disease stages prediction QNNs, QKMs
Regression Disease risk prediction Quantum linear regression
Clustering Multi-omics data analysis QK-means, quantum spectral clustering
Generative learning Synthetic medical images QNNs, QBMs, QGANs
Time series forecasting Longitudinal patient studies, drug resistance mechanisms QRC
Natural language processing Clinical notes processing QLSTM
Optimization Treatment optimization QAOA, QPSO, VQE
@@
Consider the example of dimensionality reduction, a key technique in order to compress large and complex datasets while preserving key information. As quantum computing rapidly develops, quantum algorithms for reducing the feature space size are evolving alongside it. For instance, many classical dimensionality reduction algorithms have a quantum adaptation, such as quantum principal component analysis (QPCA), quantum linear discriminant analysis (QLDA), quantum slow feature analysis (QSFA), and quantum isomap (QIsomap). QPCA speedup on quantum data is highlighted in [20]. However, the majority of the quantum algorithms for dimensionality reduction still require further algorithmic advancements and fault-tolerant hardware to achieve practical relevance.
Many other learning tasks may benefit from quantum computing. This includes supervised learning, regression analysis being one example, for instance via quantum linear regression. Likewise, unsupervised learning may benefit from approaches such as quantum k-means (QK-means) clusteringand quantum spectral clustering. Given the recent rise of generative artificial intelligence (AI) techniques, quantum generative models are receiving renewed interest, including QNNs, quantum Boltzmann machines (QBMs), and quantum generative adversarial networks (QGANs). As a final example, natural language processing is an emerging area that has recently progressed through the emergence of quantum natural language processing (QNLP).
@@ 3 Quantum computing for biomarkers
Clearly, the flow illustrates that quantum computing algorithms are only inserted at very specific points of the computational pipeline; most computational steps remain on classical systems. Optimizing the quantum part requires, for example, adapting the algorithm to both the specifics of the device and the characteristics of the data. The classical data should be embedded into a Hilbert space via data encoding techniques. Depending on the nature of the data and the specific problem at hand, it may be necessary to develop data encoding methods tailored to the application. Additionally, the quantum algorithm should account for the quantum device’s connectivity and its limitation. For instance, error mitigation and suppression protocolsshould be considered prior to running the algorithm to address potential hardware-induced errors.
@@ 3.1 Multi-dimensional data
Recent technological developments have produced an abundance of data in the domains of healthcare and life sciences. These are often correlated in intricate ways, which may necessitate a (quantum) network model approachand ML techniques such as (quantum) graph neural networks (QGNNs).
@@
While there is evidence for more data-efficient quantum machine learning algorithms, the loading of classical data into quantum computers represents a new set of challenges that exists in both current and near-term quantum hardware. How quantum computers can access data is of central importance in the field of quantum machine learning. In fact, some early quantum machine learning algorithms made assumptions around access to classical data that have been shown to lead to computational advantages over their classical counterparts. But once their data access assumptions were closely inspected and loosened, it was found that many of those computational speedups no longer hold [34]. Going deeper into the topic of quantum access to data is beyond the scope of this perspective; suffice it to say that for classical big data, data loading may erase computational quantum advantages. Different research efforts are underway, such as QRAMson the hardware side and coresetson the algorithmic side, to address this challenge. For the time being, however, small data is a more natural fit for existing and planned quantum computers.
@@ 3.1.1 Electronic health records (EHRs)
In EHRs, both the big data and small data regimes are relevant. Big data analysis occurs when considering population health problems and other trends across large cohorts. On the other hand, very often one is interested in cohorts with restrictive inclusion criteria, for instance in the context of clinical trials and rare diseases. As a result, EHR analysis looks set to gain significantly from quantum computing algorithms that allow more insights to be gained from large and small datasets.
The adoption of EHRs has been steadily increasing. They are highly promising for biomarker discovery due to the (typically) large datasets that can also capture subtle effects, relatively diverse population samples, quick data availability without the need for patient recruitment or elaborate research data collection, and longitudinal information which helps understand factors such as disease progression [35].
@@
For the scope of this paper, EHRs are assumed to contain (typically structured) information such as diagnoses, immunization, laboratory measurements, medications, procedures, and vital sign data as well as unstructured doctor notes. Also included under the EHR umbrella are social determinants of health (SDOH) and wearables. Medical images and omics, due to the explosive growth in their collection and relevance, are discussed separately. Ideal (future) EHRs would consolidate and amalgamate all such structured and unstructured health-relevant data within a single framework [35].
@@
Next to structured EHR fields, making sense of free-text doctor notes can offer a wealth of insights. Research in QNLP has been progressing quickly. The hope is that quantum algorithms can better represent linguistic information, in contrast to just replicating text patterns [36, 37]. For example, clinical notes may show that an epilepsy patient has been seizure-free [35], which would otherwise likely have to be indirectly inferred from the absence of medical codes.
SDOH are broadly defined as non-medical factors that affect health outcomes and may be grouped into five categories: economic stability, education, healthcare, neighborhood, and social context [38]. Many structured SDOH elements, such as age, as well as certain unstructured elements, such as geographic location, can often be readily extracted from EHRs. Further elements, such as social and behavioral determinants, are generally not yet captured; nevertheless, the integration between SDOH and EHRs is increasing [39]. SDOH are increasingly understood to be key predictors of morbidity and mortality in humans and even in other animals; factors such as income, social integration, and the number of adverse childhood experiences represent significant variables. Discovering and assessing the mechanisms of such biological correlates is computationally very challenging. Reasons include health selection (poor health causing individuals to be exposed to more social adversity) and the broad set of conditions that appear to be linked to SDOH, which makes selecting a single disease causation model difficult.
There are a range of challenges and opportunities associated with the secondary use of EHRs [40]. Nevertheless, the application of deep learning, ML, and AI methods to EHRs has yielded valuable medical insights over the past years, including remarkably accurate prediction of future conditions. This suggests that EHRs, including their time series data, represent a promising field for quantum machine learning. While methodologies and data quality need to further improve, deeper understanding of biomarkers is already being gained through such approaches [41, 35], particularly as EHRs are fused with other data sources [42]. The construction of such multimodal datasets may itself also benefit from AI techniques [43, 44].
@@ 3.1.2 Omics
Genomic, proteomic, and other omics data have revolutionized our understanding of complex diseases. Spurred by technological advances (including next-generation sequencing), omics often includes high-throughput empirical studies, producing raw data that necessitates multi-step computational processing and sophisticated analysis before being amenable to interpretation. In the last decade, omics have became widely used in biomedical research to study disease mechanisms, identification of biomarkers for therapeutic development, and diagnostics in the clinical setting. Although such large datasets can now be generated, biological data suffers from what is known as the “curse of dimensionality” [45], which makes it challenging to perform traditional ML, due to over-fitting, and to properly train the model. Mitigation of this is aided by state-of-the-art methods in single-cell sequencing and spatial transcriptomics, which begin to address the issue of dimensionality by interrogating individual cells rather than populations. While the number of patients will not increase with single-cell analysis, having information from individual cells allows for more detailed biomarker identification strategies, especially in diseases with heterogeneous manifestations such as cancer. However, even with single-cell analysis, limited data and too many features continue to be problems, and addressing these requires techniques such as dimensionality reduction, feature selection, and clustering. An example is biomarker discovery for the stratification of patients into different groups based on their susceptibility to a particular disease or their response to a specific drug. Often such studies are limited in the number of available samples, but owing to the large number of features the search space can be very vast. Indeed, collecting omics data from patients is challenging for many reasons, most notably the invasive nature of the medical procedures, their associated costs, and a lack of infrastructure. This is an area where quantum computing has already had demonstrable impact, and will continue to do be increasingly impactful as data availability increases and the hardware and software evolve [46, 47].
@@
With genome sequencing becoming increasingly democratized, methods such as GWAS and quantitative trait loci (QTL) mapping are often used to identify genetic biomarkers of disease states or phenotypic variation in populations. However, in order for approaches such as GWAS and QTL mapping to reach statistical significance, let alone train predictive models, huge amounts of data would be required, and this is where quantum computing can provide potential alternatives to classical algorithms [48]. In addition, GWAS studies often require methods such as PCA to identify covariates for biomarker trait regression models. In cases where the covariance matrix of interest is of a low rank, QPCAand variants thereofcan prove far more efficient than their existing classical counterparts, albeit issues of data loading and quantum hardware fault-tolerance must be overcome first (see discussion in Section 3.1).
@@
The interpretation of genetic variants in terms of biomarkers is typically limited by inherent correlations between causal and non-causal variants, requiring formal causal inference approaches. The latter are notorious for their prohibitive computational cost, except in cases with restrictive assumptions about the underlying causal structure. Quantum causal inference is an active area of research that could potentially provide more powerful causal discovery algorithms than are currently available classically. So far, quantum speedups in causal inference have been obtained only for datasets whose statistics cannot be described entirely with classical means due to the presence of quantum correlations [49].
@@
Another prevalent approach to biomarker discovery is through studies of how transcription of one gene affects the transcription of its neighboring genes, whether it be through statistical approaches such as expression quantitative trait loci (eQTLs) or gaining predictive power from gene expression patterns through machine learning. An efficient way to rapidly understand and analyze complex regulatory relationships between genes is through gene regulatory networks (GRNs). GRNs are computationally efficient since they only include pairwise interactions. Recently, a PQC model for a quantum single-cell GRN (qscGRN) was proposed to try and allow for higher-order correlations and preserving the computational efficiency of GRNs [50]. In this work, the authors analyzed single-cell RNA sequencing data from lymphoblastoid cell lines originating from two data sources [51, 52]. The authors found mixed results with the quantum models, with some correlations in accordance with previous findings, other correlations in contradiction to previous findings, and even new correlations that were not present in the baseline model. This work is an example of how novel quantum computational models and algorithms even, when co-designed with omics problems, can lead to new insights and help bypass the data uploading challenges described in earlier sections.
@@
Another avenue to understand the co-expression of genes is to implement methods to cluster gene-expression datasets. While there are many techniques for clustering gene expression data, more traditional methods often make assumptions about the data or make a priori decisions about parameters that are not necessarily sound. To circumvent this, more complex algorithms such as genetic algorithms or particle swarm optimizers (PSO) have more recently been explored for genomic clustering, with PSO showing promising results with respect to validated accuracies compared with other methods [53]. However the disadvantage of this algorithm is that it does not guarantee a convergence to a global minimum [54] and is prone to being trapped in local optima [55]. Classical techniques to overcome this drawback have been discovered, but their implementation can be computationally expensive. A potential solution to mitigate this issue is to use the quantum-behaved particle swarm optimization algorithm [54, 56] and its modified versions [57, 58, 59]. The idea is to suppose that each point-particle is now a spin-less one with quantum behavior that follows a corresponding Schrödinger equation, which can then be used to perform the particles’ position and velocity updates at each iteration. Although the references cited have implemented these quantum-behaved PSO algorithms on classical computers, to our current knowledge, they have not been discussed in the context of being implemented on a quantum computer. It appears possible that furthering the understanding of implementing quantum-behaved PSOs on quantum computers, in addition to exploring other quantum optimization techniques (not necessarily focused on clustering) such as the quantum approximate optimization algorithm (QAOA) and the variational quantum eigensolver (VQE), could lead to quantum advantages.

### v4

@@ Abstract
Biomarkers play a central role in medicine’s gradual progress towards proactive, personalized precision diagnostics and interventions. However, finding biomarkers that provide very early indicators of a change in health status, for example for multi-factorial diseases, has been challenging. Discovery of such biomarkers stands to benefit significantly from advanced information processing and means to detect complex correlations, which quantum computing offers. In this perspective paper, quantum algorithms, particularly in machine learning, are mapped to key applications in biomarker discovery. The opportunities and challenges associated with the algorithms and applications are discussed. The analysis is structured according to different data types — multi-dimensional, time series, and erroneous data — and covers key data modalities in healthcare — electronic health records (EHRs), omics, and medical images. An outlook is provided concerning open research challenges.
@@ 1 Introduction
A biomarker is a defined characteristic that is measured as an indicator of normal or pathological biological processes, or of response to an exposure or intervention [1]. The study of biomarkers arguably dates back to at least the beginning of the last century with Karl Landsteiner’s discovery of the ABO blood types in 1901 (recognized with the 1930 Nobel Prize in Physiology or Medicine). Biomarkers consist of both biological molecules [2] and other medical signs that indicate diseases [3, 4], in contrast to symptoms as well as more general medical endpoints. Biomarkers can be based on histological, molecular, physiological, and radiographic characteristics. They may also take the form of aggregate high-dimensional features.
@@
For example, clinical trials in neurology and psychiatry are time-intensive, expensive, and subjective, and have lower success rates than other indications [6]. Within neuro-psychiatry, pain represents a large unmet need with respect to the lack of reliable biomarkers. Hence, quantitative and physiological biomarkers are urgently needed for a comprehensive and reliable diagnosis and assessment of pain in healthcare and clinical trials [7]. This is illustrative of the impact better biomarkers can have across medicine.
While the beginnings of quantum mechanics date back 100 years, quantum computing is a new form of information processing, including novel hardware and software, that leverages quantum mechanical effects such as quantum entanglement, interference, and superposition. It is the only known form of computing that can provide, at least for certain problems, exponential speedups compared with classical approaches; for other problems, there may only be polynomial speedups or none at all, however [8]. Hundreds of use cases are meanwhile being explored across industry and academia [9]. Quantum computing is not just a faster way of addressing problems; it represents an entirely different way to find solutions. The benefits also go beyond speed; depending on the application and the use case, the main benefit could be accuracy [10], energy efficiency [11], or the ability to handle difficult datasets (for instance, high-dimensional or highly noisy ones) [12, 13].
The focus of this review is on (gate-based) quantum computing and its potential for computational problems in biomarker discovery. Nevertheless, it is important to note that there are further quantum technologies with medical applications and biomarker relevance, in particular quantum annealing [14] and quantum sensing [15, 16, 17]. In the rest of the paper, quantum computing and relevant quantum algorithms are first summarized in Section 2 before the application of quantum algorithms to biomarker discovery problems is described in Section 3. Open research challenges are discussed in Section 4 and conclusions are drawn in Section 5.
@@ 2 Quantum computing
In quantum computing, algorithms take advantage of Hilbert space. This is a high-dimensional vector space that represents the state of the qubits and is naturally suited to computing outcome probabilities (via an inner product). If the input of a quantum algorithm is classical data that is not represented by quantum states, data embedding is required in order to perform operations on a quantum computer. While there are various ways to embed classical data into quantum states, the number of available qubits restricts the feature space size of the data we can tackle with near-term devices. For instance, angle encoding [18] or instantaneous quantum polynomial encoding [19] require a linear scaling of the number of qubits with respect to the number of features. Alternatively, amplitude encoding provides a logarithmic scaling in the number of qubits [18]. However, without ancillary qubits, amplitude encoding results in deep circuits [20]. In practice, when the feature space exceeds the capacity of available quantum hardware, classical dimensionality reduction algorithms are applied as a pre-processing step in quantum pipelines. Still, with ongoing advancements in quantum computing, biomarker discovery and other quantum applications may eventually be performed without the need for significant (classical) pre-processing.
Quantum circuits are fundamental to the architecture of quantum computation, providing the framework through which quantum algorithms are designed and executed. At a high level, a quantum circuit is a sequence of quantum gates applied to qubits. Each gate performs a specific operation on one or more qubits. Fault-tolerant quantum algorithms typically require running deep quantum circuits on tens of thousands to millions of qubits, making them unsuitable for near-term devices. On the other hand, variational quantum algorithms (VQAs) [21] utilize shallower circuits and can produce meaningful results. This characteristic has made VQAs highly popular for near-term quantum computing applications. VQAs are hybrid quantum-classical algorithms that involve an iterative loop between quantum and classical hardware. First, quantum circuits that have parametrized gates, called the ansätze or parametrized quantum circuits (PQCs), are employed to obtain samples from a quantum computer. Then, a classical computer optimizes the parameters of the quantum circuit with respect to the loss function of the problem. Most quantum machine learning models belong in the class of VQAs and can be evaluated on quantum computers. Alternatively, quantum computers could be leveraged as sub-routines within classical algorithms. One example is the thermal state preparation during Boltzmann machine training. However, the performance of VQAs is highly sensitive to hardware noise, necessitating the use of error mitigation techniques to partially counteract noise-induced errors [22].
@@
Quantum computing is naturally suited for quantum data (which describes quantum states) [23]. However, it also holds great promise in solving problems using classical data too. In particular, there has been much progress in the last few years in understanding how quantum computers can solve computational problems that can be formulated as cryptography, machine learning (ML), optimization, or simulation (of nature) problems [24]. In this paper, we focus on analyzing classical data with the help of quantum algorithms, including multi-dimensional, time series, and erroneous data, which is discussed in detail in Section 3.
@@ 2.1 Quantum machine learning
The majority of QNNs are based on PQCs. The parameters of PQCs are optimized following the VQA paradigm. A typical QNN begins with data embedding, followed by a series of gates. One of the most general QNN models for classification is the variational quantum classifier (VQC), which is essentially a PQC. Another example of a VQC is the circuit-centric quantum classifier [26], one of the earliest proposed generic QNNs. The data re-uploading classifier is a notable QNN architecture that was shown to have universal approximation properties [27]. The data re-uploading classifier features a processing layer composed of data encoding and unitary operations with trainable parameters. This layer is repeated a number of times, each layer having independent parameters.
@@
Dimensionality reduction Covariates for biomarker trait regression models QPCA [32, 33], QLDA [34], QSFA [35], QIsomap [36]
Classification Early disease stage prediction QNNs [37, 27, 26, 38], QKMs [19, 28]
Regression Disease risk prediction Quantum linear regression [39, 40, 41, 42]
Clustering Multi-omics data analysis QK-means [43, 44], quantum spectral clustering [45]
Generative learning Synthetic medical images QNNs [46], QBMs [47, 48], QGANs [49]
Time series forecasting Longitudinal patient studies, drug resistance mechanisms QRC [50]
Natural language processing Clinical notes processing QLSTM [51]
Optimization Treatment optimization QAOA [52], QPSO [53], VQE [54]
@@
Consider the example of dimensionality reduction, a key technique in order to compress large and complex datasets while preserving key information. As quantum computing rapidly develops, quantum algorithms for reducing the feature space size are evolving alongside it. For instance, many classical dimensionality reduction algorithms have a quantum adaptation, such as quantum principal component analysis (QPCA) [32], quantum linear discriminant analysis (QLDA) [34], quantum slow feature analysis (QSFA) [35], and quantum isomap (QIsomap) [36]. QPCA speedup on quantum data is highlighted in [33]. However, the majority of the quantum algorithms for dimensionality reduction still require further algorithmic advancements and fault-tolerant hardware to achieve practical relevance.
Many other learning tasks may benefit from quantum computing. This includes supervised learning, regression analysis being one example, for instance via quantum linear regression [39, 40, 41, 42]. Likewise, unsupervised learning may benefit from approaches such as quantum k-means (QK-means) clustering [43, 44] and quantum spectral clustering [45]. Given the recent rise of generative artificial intelligence (AI) techniques, quantum generative models are receiving renewed interest, including QNNs [46], quantum Boltzmann machines (QBMs) [47, 48], and quantum generative adversarial networks (QGANs) [49]. As a final example, natural language processing is an emerging area that has recently progressed through the emergence of quantum natural language processing (QNLP) [55, 51].
@@ 3 Quantum computing for biomarkers
Clearly, the flow illustrates that quantum computing algorithms are only inserted at very specific points of the computational pipeline; most computational steps remain on classical systems. Optimizing the quantum part requires, for example, adapting the algorithm to both the specifics of the device and the characteristics of the data. The classical data should be embedded into a Hilbert space via data encoding techniques [64]. Depending on the nature of the data and the specific problem at hand, it may be necessary to develop data encoding methods tailored to the application. Additionally, the quantum algorithm should account for the quantum device’s connectivity and its limitation. For instance, error mitigation and suppression protocols [22] should be considered prior to running the algorithm to address potential hardware-induced errors.
@@ 3.1 Multi-dimensional data
Recent technological developments have produced an abundance of data in the domains of healthcare and life sciences [65]. These are often correlated in intricate ways, which may necessitate a (quantum) network model approach [66] and ML techniques such as (quantum) graph neural networks (QGNNs) [67, 38].
@@
While there is evidence for more data-efficient quantum machine learning algorithms, the loading of classical data into quantum computers represents a new set of challenges that exists in both current and near-term quantum hardware. How quantum computers can access data is of central importance in the field of quantum machine learning. In fact, some early quantum machine learning algorithms made assumptions around access to classical data that have been shown to lead to computational advantages over their classical counterparts. But once their data access assumptions were closely inspected and loosened, it was found that many of those computational speedups no longer hold [72]. Going deeper into the topic of quantum access to data is beyond the scope of this perspective; suffice it to say that for classical big data, data loading may erase computational quantum advantages. Different research efforts are underway, such as QRAMs [73] on the hardware side and coresets [74, 75] on the algorithmic side, to address this challenge. For the time being, however, small data is a more natural fit for existing and planned quantum computers.
@@ 3.1.1 Electronic health records (EHRs)
The increasing adoption of EHRs has revolutionized biomedical research. EHRs provide rapid data availability, eliminating the need for extensive recruitment or data collection. Additionally, they support longitudinal tracking of patient health, facilitating disease progression analysis. Their ability to capture data from diverse patient populations also enhances the detection of subtle health trends.
@@
For the scope of this paper, EHRs are assumed to contain (typically structured) information such as diagnoses, immunization, laboratory measurements, medications, procedures, and vital sign data as well as unstructured doctor notes. They also increasingly integrate social determinants of health (SDOH) and wearable data. Medical images and omics, due to the explosive growth in their collection and relevance, are discussed separately in this paper.
SDOH, encompassing economic stability, education, healthcare access, neighborhood environment, and social context, significantly impact health outcomes [76]. While some SDOH data (such as age and location) can be extracted from EHRs, other data (such as social and behavioral determinants) is often missing. Since SDOH are strongly correlated with morbidity and mortality, integrating them with EHRs is crucial. However, the complex interplay between health and social factors presents significant analytical challenges [77].
The application of deep learning, machine learning, and AI to EHRs has yielded valuable medical insights, including accurate prediction of future health conditions [78, 79]. AI-based approaches [80, 81] broaden the understanding of biomarkers, particularly when EHRs are fused with other data sources [82]. The construction of such multimodal datasets may itself also benefit from AI techniques [83, 84]. Therefore, the increasing integration of EHRs in biomedical research suggests that QML holds immense potential for advancing biomarker discovery. In particular, quantum computing could enhance the analysis of EHRs in small-cohort studies, such as clinical trials or rare disease research, by improving predictive accuracy from small data sets.
@@
Furthermore, quantum natural language processing (QNLP) shows promise in representing linguistic information more effectively than classical methods [55, 51]. For example, clinical notes may show that an epilepsy patient has been seizure-free [81], which would otherwise likely have to be indirectly inferred from the absence of medical codes. Therefore, QNLP methods may potentially be applied in analyzing free-text doctor’s notes, which offer valuable insights [85].
@@ 3.1.2 Omics
Genomic, proteomic, and other omics data have revolutionized our understanding of complex diseases. Spurred by technological advances (including next-generation sequencing), omics often includes high-throughput empirical studies, producing raw data that necessitates multi-step computational processing and sophisticated analysis before being amenable to interpretation. In the last decade, omics have become widely used in biomedical research to study disease mechanisms, identification of biomarkers for therapeutic development, and diagnostics in the clinical setting [86, 87]. Although such large datasets can now be generated, biological data suffers from what is known as the “curse of dimensionality” [88], which makes it challenging to perform traditional ML, due to over-fitting, and to properly train the model. Mitigation of this is aided by state-of-the-art methods in single-cell sequencing and spatial transcriptomics, which begin to address the issue of dimensionality by interrogating individual cells rather than populations. While the number of patients will not increase with single-cell analysis, having information from individual cells allows for more detailed biomarker identification strategies, especially in diseases with heterogeneous manifestations such as cancer. However, even with single-cell analysis, limited data and too many features continue to be problems, and addressing these requires techniques such as dimensionality reduction, feature selection, and clustering [89].
An example is biomarker discovery for the stratification of patients into different groups based on their susceptibility to a particular disease or their response to a specific drug. Often such studies are limited in the number of available samples, but owing to the large number of features the search space can be very vast. Indeed, collecting omics data from patients is challenging for many reasons, most notably the invasive nature of the medical procedures, their associated costs, and a lack of infrastructure. This is an area where quantum computing has already had demonstrable impact, and will continue to do be increasingly impactful as data availability increases and the hardware and software evolve [90, 91].
@@
Furthermore, with genome sequencing becoming increasingly democratized, methods such as GWAS and quantitative trait loci (QTL) mapping are often used to identify genetic biomarkers of disease states or phenotypic variation in populations [92, 86]. However, in order for approaches such as GWAS and QTL mapping to reach statistical significance, let alone train predictive models, huge amounts of data would be required, and this is where quantum computing can provide potential alternatives to classical algorithms [93]. In addition, GWAS studies often require methods such as PCA to identify covariates for biomarker trait regression models [94]. In cases where the covariance matrix of interest is of a low rank, QPCA [32] and variants thereof [95, 33] can prove far more efficient than their existing classical counterparts, albeit issues of data loading and quantum hardware fault-tolerance must be overcome first (see discussion in Section 3.1).
@@
The interpretation of genetic variants in terms of biomarkers is typically limited by inherent correlations between causal and non-causal variants [96, 97, 98], requiring formal causal inference approaches. The latter are notorious for their prohibitive computational cost [99], except in cases with restrictive assumptions about the underlying causal structure. Quantum causal inference is an active area of research that could potentially provide more powerful causal discovery algorithms than are currently available classically. So far, quantum speedups in causal inference have been obtained only for datasets whose statistics cannot be described entirely with classical means due to the presence of quantum correlations [100].
@@
Another prevalent approach to biomarker discovery with omics data is through studies of how transcription of one gene affects the transcription of its neighboring genes, whether it be through statistical approaches such as expression quantitative trait loci (eQTLs) or gaining predictive power from gene expression patterns through machine learning [101, 102]. An efficient way to rapidly understand and analyze complex regulatory relationships between genes is through gene regulatory networks (GRNs) [103, 104]. GRNs are computationally efficient since they only include pairwise interactions. Recently, a PQC model for a quantum single-cell GRN (qscGRN) was proposed to try and allow for higher-order correlations and preserving the computational efficiency of GRNs [105]. In this work, the authors analyzed single-cell RNA sequencing data from lymphoblastoid cell lines originating from two data sources [106, 107]. The authors found mixed results with the quantum models, with some correlations in accordance with previous findings, other correlations in contradiction to previous findings, and even new correlations that were not present in the baseline model. This work is an example of how novel quantum computational models and algorithms even, when co-designed with omics problems, can lead to new insights and help bypass the data uploading challenges described in earlier sections.
@@
Moreover, an additional avenue to understand the co-expression of genes is to implement methods to cluster gene-expression datasets. While there are many techniques for clustering gene expression data [108], more traditional methods often make assumptions about the data or make a priori decisions about parameters that are not necessarily sound [109]. To circumvent this, more complex algorithms such as genetic algorithms or particle swarm optimizers (PSO) have more recently been explored for genomic clustering, with PSO showing promising results with respect to validated accuracies compared with other methods [110]. However the disadvantage of this algorithm is that it does not guarantee a convergence to a global minimum [111] and is prone to being trapped in local optima [112]. Classical techniques to overcome this drawback have been discovered, but their implementation can be computationally expensive [113]. A potential solution to mitigate this issue is to use the quantum-behaved particle swarm optimization algorithm [111, 114] and its modified versions [115, 116, 117]. The idea is to suppose that each point-particle is now a spin-less one with quantum behavior that follows a corresponding Schrödinger equation, which can then be used to perform the particles’ position and velocity updates at each iteration. Although the references cited have implemented these quantum-behaved PSO algorithms on classical computers, to our current knowledge, they have not been discussed in the context of being implemented on a quantum computer. It appears possible that furthering the understanding of implementing quantum-behaved PSOs on quantum computers, in addition to exploring other quantum optimization techniques (not necessarily focused on clustering) such as the quantum approximate optimization algorithm (QAOA) and the variational quantum eigensolver (VQE), could lead to quantum advantages.